Mary and Mark are saving to buy new bikes. They both start with $20. If Mary saves $15 a week and Mark saves $20 a week, how much will each have saved in week 6? Make a table and graph to solve.

Asap

Answer by 229388(1)



Let Mary = a
Let Mark = b
Let w = number of weeks

We know that: a=20+15w
b=20+20w

In 6 weeks, just plug in w=6.

A = 20 + 15*6 = 110
B = 20 + 20*6 = 140

Mary: 110 dollars
Mark: 140 dollars